OpenStudy (anonymous):

1/2 log[5]1-2 log[5]5

OpenStudy (michele_laino):

hint: \[{\log _5}1 = 0\]

OpenStudy (anonymous):

This is how far my daughter's gotten. Not sure if she's on the right track or not.

OpenStudy (michele_laino):

Here are the right steps: \[\frac{1}{2}{\log _5}1 - 2{\log _5}5 = \frac{1}{2} \cdot 0 - 2 \cdot 1 = 0 - 2 = - 2\]